OpenStudy (elleblythe):

How to solve for sec(-pi/9)sin(-23pi/6)cos(pi/9) using the values on the unit circle and not decimal values??

OpenStudy (aum):

sec(-x) = sec(x). True for cosine too. cos(-x) = cos(x). sec(-pi/9) = sec(pi/9) = 1 / cos(pi/9) sec(-pi/9) * sin(-23pi/6) * cos(pi/9) = ( 1 / cos(pi/9) ) * sin(-23pi/6) * cos(pi/9) = sin(-23pi/6) Keep adding as many multiples of 2pi and sine value will not change. Therefore, -23pi/6 + 2pi = pi( -23/6 + 2) = pi( (-23+12)/6 ) = pi(-11/6) Add another 2pi pi(-11/6) + pi(2) = pi(-11/6 + 2) = pi( (-11 + 12) / 6 ) = pi/6 sin(-23pi/6) = sin(-11pi/6) = sin(pi/6) = 1/2

OpenStudy (elleblythe):

@aum what happens to the cos(pi/9) though?

OpenStudy (aum):

cos(pi/9) cancels out sec(pi/9). cos(pi/9) times sec(pi/9) is 1. They are reciprocals of each other. Since sec(pi/9) = 1 / cos(pi/9), cos(pi/9) x sec(pi/9) = cos(pi/9) x 1 / cos(pi/9) = 1